What would be the new volume if the temperature on 450 mL of gas is changed from 45oC to -5oC?




1
Expert's answer
2021-01-22T06:20:51-0500

T1 = 450C = 45 + 273.15 = 318.15 K

T2 = -50C = -5 + 273.15 = 268.15 K

V1 = 450 mL

V2 = V1 T2 / T1 = (450 x 268.15) / 318.15 =

= 379.3 mL